A hang glider dropped his cell phone from a height of 350 feet. How many seconds did it take for the cell phone to reach the ground?
Answer by Alan3354(69443)

A hang glider dropped his cell phone from a height of 350 feet. How many seconds did it take for the cell phone to reach the ground?
-----------
s = at^2/2
350 = 32*t^2/2 = 16t^2
t^2 = 350/16
t =~ 4.677 seconds
============
Was that hard?